OpenStudy (anonymous):

Convert r=5/(1 +cos theta) to a rectangular equation.

OpenStudy (anonymous):

\[r={5\over1+\cos\theta}\\ r(1+\cos\theta)=5\\ r+r\cos\theta=5 \] we have the relation: \[r=\sqrt{x^2+y^2}\\ x=r\cos\theta\qquad y=r\sin\theta\] plug them in

OpenStudy (anonymous):

\[\sqrt{x^2+y^2}+x=5\\ \sqrt{x^2+y^2}=5-x \] square both sides and simplify